how much work is required to move a charge of 4nC from a point 2m away to a point 0.5m away from a point charge of 60nC?what is the potential difference between these points?


1
Expert's answer
2021-04-19T17:43:01-0400

The potential at point 2m away is-


"V_1=\dfrac{1}{4\pi \epsilon_o}\dfrac{q}{r}=9\times 10^9\times \dfrac{60\times 10^{-9}}{2}=270 {\text{volts}}"

The potential at point 0.5m away is-


"V_2=\dfrac{1}{4\pi \epsilon_o}\dfrac{q}{r}=9\times 10^9\times \dfrac{60\times 10^{-9}}{0.5}=1080 {\text{volts}}"



Then The potential difference between these points is-

"V=V_2-V_1=1080-270=810 \text{volts}"



Also, Work done required "=4\times 10^{-9}\times V=4\times 10^{-9}\times 810=3.24\times 10^{-6} J"